# Manifold Learning by Mixture Models of VAEs for Inverse Problems

### v2

@@ Appendix B Error Bound for the Approximation of $\beta_{ik}$ by $\tilde{\beta}_{ik}$
It is well-known that the difference between evidence and ELBO can be expressed in terms of the Kullback-Leibler divergence, see [62, Sec 2.2]. In the special case that $E\circ D=\mathrm{Id}$, which is relevant in this paper, this estimate can be simplified by the following lemma. To this end, denote by
@@
Proof.
By [62, eqt (2.8)], it holds that
@@
Using $\xi\sim\mathcal{N}(0,\sigma_{z}^{2}I_{d})$ and $\eta=\mathcal{N}(0,\sigma_{x}^{2}I_{n})$, we arrive at the assertion. ∎
@@
Proof.
Due to Lemma 12, we have that
@@
which concludes the proof. ∎
Lemma 14.
@@
Proof.
We estimate the density $p_{E_{\#}\mathcal{N}(x,\sigma_{x}^{2}I_{n}))}(z)$ from below. By [6, Lemma 4], we have that
@@
which proves the claim. ∎
@@
Acknowledgements
This material is based upon work supported by the Air Force Office of Scientific Research under award number FA8655-20-1-7027. We acknowledge the support of Fondazione Compagnia di San Paolo. Co-funded by the European Union (ERC, SAMPDE, 101041040). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. GSA, MS and SS are members of the “Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni”, of the “Istituto Nazionale di Alta Matematica”. JH acknowledges funding by the German Research Foundation (DFG) within the project STE 571/16-1 and by the EPSRC programme grant “The Mathematics of Deep Learning” with reference EP/V026259/1.

### v3

@@ Appendix B Error Bound for the Approximation of $\beta_{ik}$ by $\tilde{\beta}_{ik}$
It is well-known that the difference between evidence and ELBO can be expressed in terms of the Kullback-Leibler divergence (see Kingma and Welling, 2019, Sec 2.2). In the special case that $E\circ D=\mathrm{Id}$, which is relevant in this paper, this estimate can be simplified by the following lemma. To this end, denote by
@@
Proof  By Equation (2.8) in Kingma and Welling (2019), it holds that
@@
Using $\xi\sim\mathcal{N}(0,\sigma_{z}^{2}I_{d})$ and $\eta=\mathcal{N}(0,\sigma_{x}^{2}I_{n})$, we arrive at the assertion.
@@
Proof  Due to Lemma 11, we have that
@@
which concludes the proof.
Lemma 13
@@
Proof  We estimate the density $p_{E_{\#}\mathcal{N}(x,\sigma_{x}^{2}I_{n}))}(z)$ from below. By (Altekrüger et al., 2023, Lemma 4), we have that
@@
which proves the claim.